OpenStudy (anonymous):

te^(-t) sin2t can you solve this please using convolution transform in laplace

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/DE/ConvolutionIntegrals.aspx there's a whole section devoted to laplace transform

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

i am working on it .

OpenStudy (anonymous):

\(te^{-t}\sin 2t\) or \(te^{-t}\sin^2t\)?

OpenStudy (anonymous):

the first one .. pls answer tnx

OpenStudy (anonymous):

Do you want its Laplace transform or what?

OpenStudy (anonymous):

convolution sir or transform of integral

OpenStudy (anonymous):

You haven't given anything to *solve*. What are you trying to determine?

OpenStudy (anonymous):

te^-t sin2t

OpenStudy (anonymous):

here we go !!!

OpenStudy (anonymous):

sir sorry for over demanding ... did you use integration by parts?

OpenStudy (anonymous):

no i just used defination of convolution theorem .

OpenStudy (anonymous):

you need to find the inverse laplace transform of what @sami-21 got.

OpenStudy (anonymous):

the convolution formula is this sir \[\int\limits_{0}^{t}\] f(t) g(t-T)dt

OpenStudy (anonymous):

i just need the answer of the convolution sir b4 transforming it on laplace

OpenStudy (anonymous):

yes you use it for inverse process . if you have the inverse and you want to find the Laplace then go for that way . but when you need in the forward way simply find the laplace of f(t) and g(t) Multiply thm .

OpenStudy (anonymous):

ahhh sir this is last can u answer in inverse way tnx

OpenStudy (anonymous):

convolution first sir then you inverse it

OpenStudy (anonymous):

Oh. You want \(te^{-t}*\sin2t\)?

OpenStudy (anonymous):

yes sir convolution first then you inverse laplace it tnx

OpenStudy (anonymous):

\[\Large L (f*g)t=F(s)G(s)\] here \[\Large f(t)=te^{-t}\] its Laplace is \[\Large F(s)=\frac{1}{(s+1)^2}\] \[\Large g(t)=\sin(2t)\] \[\Large G(s)=\frac{2}{s^2+4}\] just use the above mentioned convolution theorem to get the required result just Multiply F(s)G(s) this is the answer .

OpenStudy (anonymous):

Now I'm really confused. I don't understand the question.

OpenStudy (anonymous):

$$\begin{align*} \mathcal{L}\{te^{-t}*\sin2t\}&=\mathcal{L}\{te^{-t}\}\mathcal{L}\{\sin2t\}\\ &=\frac1{(s+1)^2}\cdot\frac2{s^2+4}\\ &=\frac2{(s+1)^2(s^2+4)}\\ \end{align*}$$

OpenStudy (anonymous):

Use partial fraction decomposition to break into a sum (good luck solving the ugly system of linear equations) and take the Laplace transform of each of those parts.

OpenStudy (anonymous):

sir i just knew that in using convolution u don't need to transform in f(s) right .. u will answer it by using only f(t), after u answer it by using this formula\[\int\limits_{0}^{t}f(t)g(t-T)dt \] and using integration by parts on it , then you can inverse transform i.. thats my point

OpenStudy (anonymous):

That is the definition of convolution. You can evaluate that integral if you want.

OpenStudy (anonymous):

pls integrate it sir by using by parts but first you follow the formula in convolution tnx

OpenStudy (anonymous):

If you didn't want to use Laplace transforms why did you said you did? $$\begin{align*}te^{-t}*\sin2t&=\int_0^t\tau e^{-(t-\tau)}\sin2\tau\,\mathrm{d}\tau\\&=\int_0^te^{-t}\tau e^\tau\sin2\tau\,\mathrm{d}\tau\\&=e^{-t}\int_0^t\tau e^\tau\sin2\tau\,\mathrm{d}\tau\end{align*}$$ Now, let \(u=\tau\) and \(\mathrm{d}v=e^\tau\sin2\tau\,\mathrm{d}\tau\); thus \(\mathrm{d}u=\mathrm{d}\tau\) and \(v=\int e^\tau\sin2\tau\,\mathrm{d}\tau=e^\tau\sin2\tau+\frac12\int e^\tau\cos2\tau\,\mathrm{d}\tau=\dots=\frac15e^\tau(\sin2\tau-2\cos2\tau)\). I leave the rest to you!

OpenStudy (anonymous):

tnhk you so much